# Compact Language Models via Pruning and Knowledge Distillation

### v1

@@ Abstract
Large language models (LLMs) targeting different deployment scales and sizes are currently produced by training each variant from scratch; this is extremely compute-intensive. In this paper, we investigate if pruning an existing LLM and then re-training it with a fraction (<3%) of the original training data can be a suitable alternative to repeated, full retraining. To this end, we develop a set of practical and effective compression best practices for LLMs that combine depth, width, attention and MLP pruning with knowledge distillation-based retraining; we arrive at these best practices through a detailed empirical exploration of pruning strategies for each axis, methods to combine axes, distillation strategies, and search techniques for arriving at optimal compressed architectures. We use this guide to compress the Nemotron-4 family of LLMs by a factor of 2-4$\times$, and compare their performance to similarly-sized models on a variety of language modeling tasks. Deriving 8B and 4B models from an already pretrained 15B model using our approach requires up to 40$\times$ fewer training tokens per model compared to training from scratch; this results in compute cost savings of 1.8$\times$ for training the full model family (15B, 8B, and 4B). Minitron models exhibit up to a 16% improvement in MMLU scores compared to training from scratch, perform comparably to other community models such as Mistral 7B, Gemma 7B and Llama-3 8B, and outperform state-of-the-art compression techniques from the literature. We have open-sourced Minitron model weights on Huggingface  [^1] https://huggingface.co/nvidia, with corresponding supplementary material including example code available on GitHub  [^2] https://github.com/NVlabs/Minitron.
@@ 2 Pruning Methodology
Figure 2: High-level overview of our proposed iterative pruning and distillation approach to train a family of smaller LLMs.
@@ 2.1 Background and Notation
where $\mathbf{X}$ denotes the input, and $\boldsymbol{W_{1}}$ and $\boldsymbol{W_{2}}$ are the two associated weight matrices in the MLP layer. $\boldsymbol{W}_{1},\boldsymbol{W}_{2}\in\mathbb{R}^{d_{hidden}\times d_{model}}$ where $d_{model}$ and $d_{hidden}$ are the embedding and MLP hidden dimensions, respectively. $\delta(\cdot)$ refers to the non-linear activation function.
@@
$\operatorname{MHA}(\mathbf{X})$ $=\operatorname{Concat}(\text{head}_{1},...\text{head}_{L})\cdot\boldsymbol{W}^{O},$
$\text{head}_{i}=\operatorname{Attn}(\mathbf{X}\boldsymbol{W}^{Q,i},\mathbf{X}\boldsymbol{W}^{K,i},\mathbf{X}\boldsymbol{W}^{V,i}),$
@@
Finally, the Layer Normalization operation (LayerNorm) [5] on an input $\mathbf{X}$ is defined as follows,
@@ 2.2 Importance Analysis
Here, $\boldsymbol{W}_{1}^{i}$ refers to the $i^{\text{th}}$ row of the weight matrix $\boldsymbol{W_{1}}$. $\sum_{\mathbf{B,S}}$ refers to aggregation along the batch and sequence dimensions. We observe from our experiments that performing a simple summation here is not always optimal. To this end, we perform a detailed evaluation of various aggregation functions along each of these dimensions and their corresponding performance in Table 15(Appendix). Specifically, for a sequence of scores $\mathbf{S}$, we try three functions: (1) mean: $\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}_{i}$, (2) L2 norm: $\sqrt{\sum_{i=1}^{n}\mathbf{S}_{i}^{2}}$, and (3) variance: $\frac{1}{n}\sum_{i=1}^{n}(\mathbf{S}_{i}-\bar{\mathbf{S}})^{2}$. Layer-wise scores are then summed up to obtain network-wide importance scores for each axis.
@@ 2.3 Obtaining a Pruned Model
Figure 3: Overview of our neural architecture search algorithm. RT refers to retraining.
@@ 3 Retraining
where $\tau$ is the softmax temperature and ${|V|}$ is the vocabulary size. Logit-based KD loss across the sequence of all output tokens is represented as
@@
For distillation, we explore various loss functions, and several combinations of intermediate states and mappings across the Transformer model as the loss components, along with their respective trade-offs. This is illustrated in Figure 4. The intermediate state-based KD loss across a sequence of Transformer-specific hidden states is represented as
@@ 4 Experiments and Results
Data and Training Hyperparameters: we use the Nemotron-4 curated 8 trillion token (8T) base pretraining dataset and the continued training dataset (CT)  [42]. We use the 8T training blend for all our ablations and use a combination of both data blends to retrain our final models. Unless otherwise specified, we use 1.8 billion tokens (400 steps) for lightweight retraining. The calibration dataset $D$ used for importance estimation consists of 1024 samples drawn randomly from the full dataset. We use the same optimizer settings and data split as [42] with cosine LR decay schedule from $2^{-4}$ to $4.5^{-7}$.
Models Benchmark Metric Llama-3 Llama-2 Mistral Gemma Nemotron-4 Nemotron-3 Minitron # Parameters 8B 6.7B 7.3B 8.5B 15.6B 8.5B 8.3B # Non-Emb. Params 5.9B 6.4B 7B 7.7B 12.5B 6.4B 6.2B # Training Tokens >15T 2T 8T 6T 8T 3.8T 94B Knowledge,Logic winogrande (5) acc 78 74 78.5 78 83.6 75.9 79.0 arc_challenge (25) acc_norm 58 53 60.3 61 58.8 52.8 52.6 MMLU(5) acc 65 46 64.1 64 66.6 54.7 63.8 hellaswag(10) acc_norm 82 79 83.2 82 84.6 78.5 80.7 gsm8k(5) acc 50 14 37 50 48.5 24.0 51.3 truthfulqa(0) mc2 44 39 42.6 45 40.7 36.5 42.6 XLSum en (20)(3) rougeL 31 31 4.80 17 32 30.9 31.2 Coding MBPP(0) pass@1 42 20 38.8 39 38 27.04 35.2 humaneval (n=20)(0) pass@1 28 12 28.7 32 35.4 20.7 31.6
@@ 4.1 Main Pruning Results
We arrive at this list through a detailed set of ablations and experiments, and each point is backed by empirical evidence, as we demonstrate in the rest of this section and the Appendix. We use this list to obtain our Minitron pruned and retrained models, whose performance is shown in Tables 3 and 4. Here, we compare the performance of our pruned models to multiple baselines: (1) the original Nemotron-4 15B model, (2) the previous generation Nemotron-3 8B model, and (3) a set of similarly-sized community models, all trained from scratch with trillions of tokens. Evaluation is performed on the downstream tasks described earlier in this Section. In both tables, we list the number of full and non-embedding parameters, along with the number of training tokens used to arrive at the model.
We further compare the Minitron models to state-of-the-art depth and width-pruned baselines in Table  5; namely, LLM-Pruner [33], SliceGPT [4], LaCo [53], ShortGPT [34], and Sheared LLaMa [51]. Table 2 lists the architecture details of the Nemotron and Minitron models shown in Tables 3 and 4. In the following subsections, we will go into more detail on how we arrived at the Minitron pruned models.
@@
to better understand how Minitron models perform after supervised fine-tuning (SFT), we perform SFT on Minitron 4B using instruction-tuning data used for Nemotron-4 340B [38] to create Minitron 4B-instruct, and evaluate it on various tasks, including instruction-following and roleplay (IFEval and MT-Bench), RAG QA (ChatRAG-Bench), and function calling (BFCL). The results for this experiment are shown in Tables 7 to 9. Tables 7 to 9 demonstrate that Minitron 4B-instruct has strong instruction-following, roleplay and RAG capabilities, beating similarly sized models across all tasks. On function calling (Table  9), Minitron 4B-instruct outperforms Gemma-2B-IT and even Llama-3-8B-instruct.
Best Practice #1: in summary, Tables 39 provide strong empirical evidence to support the claim that training one single big model, and obtaining smaller ones from it through pruning + retraining achieves higher accuracy and is extremely cost/compute-efficient when compared to training them from scratch. Further, our efficient retraining strategy also eliminates the need to curate trillions of tokens of data.
@@
Models Benchmark Metric Phi-2 Gemma Gemma2* Qwen2* MiniCPM* Minitron # Parameters 2.7B 2.5B 2.6B 1.5B 2.7B 4.2B # Non-Emb. Params 2.5B 2B 2B 1.3B 2.4B 2.6B # Training Tokens 1.4T 3T 2T 7T 1.1T 94B Knowledge, Logic winogrande (5) acc 74 67 70.9 66.2 - 74 arc_challenge (25) acc_norm 61 48 55.4 43.9 - 50.9 MMLU(5) acc 57.5 42 51.3 56.5 53.5 58.6 hellaswag(10) acc_norm 75.2 72 73 66.6 68.3 75 gsm8k(5) acc 55 18 23.9 58.5 53.8 24.1 truthfulqa(0) mc2 44 33 - 45.9 - 42.9 XLSum en (20)(3) rougeL 1 11 - - - 29.5 Coding MBPP(0) pass@1 47 29 29.6 37.4 - 28.2 humaneval (n=20)(0) pass@1 50 24 17.7 31.1 - 23.3
@@ 4.2 Obtaining the Best Pruned Model
Best Aggregation Metric (Best Practice #2): we start by exploring the best aggregation metric for use with our activation-based pruning criteria (see Section 2.2 for more details). Table 15 (Appendix) shows how zero-shot LM loss and Wikitext2 perplexity [35] vary w.r.t different intra-batch and sequence aggregation functions. Here, the Nemotron-4 15B model is pruned to the Nemotron-3 8B architecture with no retraining. We notice that there is significant variation in zero-shot performance based on the aggregation metric, indicating the importance of selecting the right one. Both (batch=L2, seq=mean) and (mean, mean) perform well; in the remainder of the paper, we use (l2, mean) primarily due to its slightly better performance on the 8T dataset. To further evaluate if these relative rankings hold after retraining, we perform a related experiment: we prune the same 15B model to 8B using: (1) the best ((L2, mean) metric, and (2) a poorly performing (L2, L2) metric, and perform retraining on both for 400 steps ($\sim$1.8B tokens) The results of this experiment are shown in Figure 5 (Appendix). From the Figure, we conclude that these rankings continue to hold post-retraining.
@@
Minitron 4B-instruct 64.23
Gemma-2B-IT 47.00
Llama-3-8B-instruct 58.78
Table 9: Evaluation results on BFCL.
@@
T=1 5.43 1.92
T=2 5.55 1.92
T=4 5.24 1.92
Table 10: Minitron LM loss values for EMB pruning across iterations, before and after retraining.
Combining Depth and Width (Best Practice #4): we perform a simple experiment to compare the efficacy of width vs. depth pruning. Using the PPL and BI metrics defined in Section 2.2, we remove the 16 least important layers from the Nemotron 15B model based on both metrics to arrive at two variants of depth pruned models. We also perform neuron, head and embedding channel pruning to target the Nemotron-3 8B model and arrive at the width pruned variant. Finally, we combine depth (remove 4 least important layers) and width pruning to arrive at the fourth variant. We report the results of this experiment in Table 11. We notice that even though the depth-width pruned variant has a lower loss post-pruning, we see the results flip around 200 steps of retraining (0.8B tokens); Table 1 and Figure 6 (Appendix) further illustrate this.
@@
Table 11: Depth vs width pruning. LM loss comparison post retraining with 1.8B tokens.
@@ 4.3 Retraining and Search
Distillation vs. Conventional Training (Best Practice #5): in this experiment, we train a 4B parameter model and compare: (1) train with random initialization (4B-Random-Init); prune 15B to 4B, then (2) retrain with conventional training (4B-Pruned), and (3) retrain with distillation using the 15B model as the teacher (4B-Pruned-Distill). Since distillation adds training overheads (additional forward pass on the teacher model), we compare approaches under iso-compute settings. The results for this experiment are shown in Table 12. Here, we observe a significant improvement in MMLU for (3) while both (1) and (2) score randomly. On HellaSwag, (3) > (2) > (1). This clearly demonstrates the superiority of using distillation over conventional training after pruning.
@@
Table 12: Accuracy comparison for 4B using the 8T blend. Indicates settings with iso-compute.
Choice of Loss Function (Best Practice #5): we experiment with Kullback-Leibler divergence (KLD), MSE, cosine similarity and reverse KLD (R-KLD) to compute $L_{logits}$. Recent work has shown R-KLD [15, 27] to be a better fit than KLD in the SFT/instruction-following setting, and Agarwal et al. [2] claim the choice of loss is task-dependent. We observe from Table 18 and 18 (Appendix) that KLD is the best choice for pruned base model training.
Choice of Losses (Best Practices #6 and #7): typically, a weighted combination of $L_{CLM}$ and $L_{logits}$ is used for distillation. We find that using $L_{logits}$ alone results in the best performance as shown in Table 18 (Appendix). For $L_{is}=L_{emb}+L_{att}+L_{i}+L_{o}$ , we make several observations similar to Lu et al. [31]; these are listed in Appendix A.3 and in Table 19. Most notably, we observe no improvements from using $L_{is}$ when retraining models that don’t prune the depth axis significantly, such as Minitron 8B and Minitron 4B and hence use $L_{logits}$ alone in such cases (see Table 16 in the Appendix).
One-shot vs Iterative Pruning and Distillation Across Model Sizes (Best Practice #8): compressing Nemotron-4 15B to Minitron 4B requires an aggressive 73.3% reduction of original model weights. We hypothesize that aggressive one-shot pruning loses out on important capabilities of the base LLM. We thus explore a simple iterative two-step pruning and retraining strategy where we first prune and retrain Nemotron-4 15B to create Minitron 8B ($\sim$46reduction) and further prune and retrain the latter to Minitron 4B ($\sim$50reduction). Table 12 shows the comparison between single-shot and iterative pruning, and demonstrates that iterative achieves a 12% improvement in the MMLU scores compared to the one-shot strategy. During the final retraining step, we observe that using Nemotron-4 15B as the teacher achieves superior results compared to using Minitron 8B.
Search with Retraining (Best Practice #9): For lightweight neural architecture search, we use the search spaces defined in Table 13 for Minitron 8B and 4B.
@@
We further specify a target parameter range of 8 and 4 billion parameters for the respective models, with a tolerance of 5%. With these settings, we obtain 15 and 18 feasible candidates for the 8B and 4B parameter targets, respectively. The architecture configurations for these candidates are provided in Table 20 (Appendix). As described in Section 2.3, we perform lightweight retraining of all feasible candidates. Figure 10 illustrates how validation loss changes for the 8B candidates as training progresses. We notice that relative rankings undergo significant changes up to $\sim$ 300 steps, and then stabilize.
@@
Strategy Tokens MMLU HellaSwag PIQA HumanEval
Phase1 + Phase2 113B 54.7 80.3 77.2 25.6
Phase2 only 94B 61.9 80.1 76.7 30.5

### v2

@@ Abstract
Large language models (LLMs) targeting different deployment scales and sizes are currently produced by training each variant from scratch; this is extremely compute-intensive. In this paper, we investigate if pruning an existing LLM and then re-training it with a fraction (<3%) of the original training data can be a suitable alternative to repeated, full retraining. To this end, we develop a set of practical and effective compression best practices for LLMs that combine depth, width, attention and MLP pruning with knowledge distillation-based retraining; we arrive at these best practices through a detailed empirical exploration of pruning strategies for each axis, methods to combine axes, distillation strategies, and search techniques for arriving at optimal compressed architectures. We use this guide to compress the Nemotron-4 family of LLMs by a factor of 2-4$\times$, and compare their performance to similarly-sized models on a variety of language modeling tasks. Deriving 8B and 4B models from an already pretrained 15B model using our approach requires up to 40$\times$ fewer training tokens per model compared to training from scratch; this results in compute cost savings of 1.8$\times$ for training the full model family (15B, 8B, and 4B). Minitron models exhibit up to a 16% improvement in MMLU scores compared to training from scratch, perform comparably to other community models such as Mistral 7B, Gemma 7B and Llama-3 8B, and outperform state-of-the-art compression techniques from the literature. We have open-sourced Minitron model weights on Huggingface  [^1] Minitron Collection on HuggingFace, with corresponding supplementary material including example code available on GitHub  [^2] https://github.com/NVlabs/Minitron.
@@ 2 Pruning Methodology
Figure 2: High-level overview of our proposed iterative pruning and distillation approach to train a family of smaller LLMs. On a pretrained LLM, we first evaluate importance of neurons, rank them, trim the least important neurons and distill the knowledge from the original LLM to the pruned model. The original model is replaced with the distilled model for the next iteration of compression.
@@ 2.1 Background and Notation
where $\mathbf{X}$ denotes the input, and $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ are the two associated weight matrices in the MLP layer. $\boldsymbol{W}_{1},\boldsymbol{W}_{2}\in\mathbb{R}^{d_{hidden}\times d_{model}}$, where $d_{model}$ and $d_{hidden}$ are the embedding and MLP hidden dimensions, respectively. $\delta(\cdot)$ refers to the non-linear activation function.
@@
$\operatorname{MHA}(\mathbf{X})=\operatorname{Concat}(\text{head}_{1},...\text{head}_{L})\cdot\boldsymbol{W}^{O}$
$\text{head}_{i}=\operatorname{Attn}(\mathbf{X}\boldsymbol{W}^{Q,i},\mathbf{X}\boldsymbol{W}^{K,i},\mathbf{X}\boldsymbol{W}^{V,i})$
@@
Finally, the Layer Normalization operation (LayerNorm) [5] on an input $\mathbf{X}$ is defined as follows:
@@ 2.2 Importance Analysis
Here, $\boldsymbol{W}_{1}^{i}$ refers to the $i^{\text{th}}$ row of the weight matrix $\boldsymbol{W_{1}}$. $\sum_{\mathbf{B,S}}$ refers to aggregation along the batch and sequence dimensions. We observe from our experiments that performing a simple summation here is not always optimal. To this end, we perform a detailed evaluation of various aggregation functions along each of these dimensions and their corresponding performance in Table 13. Specifically, for a sequence of scores $\mathbf{S}$, we try three functions: (1) mean(abs): $\frac{1}{n}\sum_{i=1}^{n}|\mathbf{S}_{i}|$ (hereafter referred to as just mean), (2) L2 norm: $\sqrt{\sum_{i=1}^{n}\mathbf{S}_{i}^{2}}$, and (3) variance: $\frac{1}{n}\sum_{i=1}^{n}(\mathbf{S}_{i}-\bar{\mathbf{S}})^{2}$. Layer-wise scores are then summed up to obtain network-wide importance scores for each axis.
@@ 2.3 Obtaining a Pruned Model
Figure 3: Overview of our neural architecture search algorithm. We perform a search on multiple axes: number of layers, attention head count, MLP and embedding dimensions to arrive at a set of feasible architectures meeting the parameter budget. RT refers to retraining.
@@ 3 Retraining
where $\tau$ is the softmax temperature and ${|V|}$ is the vocabulary size. Logit-based KD loss across the sequence of all output tokens is represented as:
@@
For distillation, we explore various loss functions, and several combinations of intermediate states and mappings across the Transformer model as the loss components, along with their respective trade-offs. This is illustrated in Figure 4. The intermediate state-based KD loss across a sequence of Transformer-specific hidden states is represented as:
@@ 4 Experiments and Results
Data and Training Hyperparameters: we use the Nemotron-4 curated 8 trillion token (8T) base pretraining dataset and the continued training dataset (CT)  [42, 44, 43]. We use the 8T training blend for all our ablations and use a combination of both data blends to retrain our final models. Unless otherwise specified, we use 1.8 billion tokens (400 steps) for lightweight retraining. The calibration dataset $D$ used for importance estimation consists of 1024 samples drawn randomly from the full dataset. We use the same optimizer settings and data split as [43] with cosine LR decay schedule from $2^{-4}$ to $4.5^{-7}$.
Models Benchmark Metric Llama-3 Llama-2 Mistral Gemma Nemotron-4 Nemotron-3 Minitron # Parameters 8B 6.7B 7.3B 8.5B 15.6B 8.5B 8.3B # Non-Emb. Params 5.9B 6.4B 7B 7.7B 12.5B 6.4B 6.2B # Training Tokens >15T 2T 8T 6T 8T 3.8T 94B Knowledge,Logic winogrande (5) acc 77.6 74 78.5 78 83.6 75.9 79.0 arc_challenge (25) acc_norm 57.8 53 60.3 61 58.8 52.8 52.6 MMLU(5) acc 65.3 46 64.1 64 66.6 54.7 63.8 hellaswag(10) acc_norm 82.1 79 83.2 82 84.6 78.5 80.7 gsm8k(5) acc 50.3 14 37 50 48.5 24.0 51.3 truthfulqa(0) mc2 43.9 39 42.6 45 40.7 36.5 42.6 XLSum en (20)(3) rougeL 30.9 31 4.80 17 32 30.9 31.2 Coding MBPP(0) pass@1 42.4 20 38.8 39 38 27.04 35.2 humaneval (n=20)(0) pass@1 28.1 12 28.7 32 35.4 20.7 31.6
@@
Models Benchmark Metric Phi-2 Gemma Gemma2* Qwen2* MiniCPM* Minitron # Parameters 2.7B 2.5B 2.6B 1.5B 2.7B 4.2B # Non-Emb. Params 2.5B 2B 2B 1.3B 2.4B 2.6B # Training Tokens 1.4T 3T 2T 7T 1.1T 94B Knowledge, Logic winogrande (5) acc 74 67 70.9 66.2 - 74.0 arc_challenge (25) acc_norm 61 48 55.4 43.9 - 50.9 MMLU(5) acc 57.5 42 51.3 56.5 53.5 58.6 hellaswag(10) acc_norm 75.2 72 73.0 66.6 68.3 75.0 gsm8k(5) acc 55 18 23.9 58.5 53.8 24.1 truthfulqa(0) mc2 44 33 - 45.9 - 42.9 XLSum en (20)(3) rougeL 1 11 - - - 29.5 Coding MBPP(0) pass@1 47 29 29.6 37.4 - 28.2 humaneval (n=20)(0) pass@1 50 24 17.7 31.1 - 23.3
@@ 4.1 Main Pruning Results
We arrive at this list through a detailed set of ablations and experiments, and each point is backed by empirical evidence, as we demonstrate in the rest of this section and the Appendix. We use this list to obtain our Minitron pruned and retrained models, whose performance is shown in Tables 2 and 3. Here, we compare the performance of our pruned models to multiple baselines: (1) the original Nemotron-4 15B model, (2) the previous generation Nemotron-3 8B model, and (3) a set of similarly-sized community models, all trained from scratch with trillions of tokens. Evaluation is performed on the downstream tasks described earlier in this Section. In both tables, we list the number of full and non-embedding parameters, along with the number of training tokens used to arrive at the model.
We further compare the Minitron models to state-of-the-art depth and width-pruned baselines in Table 4; namely, LLM-Pruner [33], SliceGPT [4], LaCo [55], ShortGPT [34], and Sheared LLaMa [53]. Table 5 lists the architecture details of the Nemotron and Minitron models shown in Tables 2 and 3. In the following subsections, we will go into more detail on how we arrived at the Minitron pruned models.
@@
to better understand how Minitron models perform after supervised fine-tuning (SFT), we perform SFT on Minitron 4B using instruction-tuning data used for Nemotron-4 340B [38] to create Minitron 4B-instruct, and evaluate it on various tasks, including instruction-following and roleplay (IFEval and MT-Bench), RAG QA (ChatRAG-Bench), and function calling (BFCL). The results for this experiment are shown in Tables 7 to 9. Tables 7 to 9 demonstrate that Minitron 4B-instruct has strong instruction-following, roleplay and RAG capabilities, beating similarly sized models across all tasks. On function calling (Table 9), Minitron 4B-instruct outperforms Gemma-2B-IT and even Llama-3-8B-instruct.
Best Practice #1: in summary, Tables 2 - 9 provide strong empirical evidence to support the claim that training one single big model, and obtaining smaller ones from it through pruning + retraining achieves higher accuracy and is extremely cost/compute-efficient when compared to training them from scratch. Further, our efficient retraining strategy also eliminates the need to curate trillions of tokens of data.
@@ 4.2 Obtaining the Best Pruned Model
Best Aggregation Metric (Best Practice #2): we start by exploring the best aggregation metric for use with our activation-based pruning criteria (see Section 2.2 for more details). Table 13 shows how zero-shot LM loss and Wikitext2 perplexity [35] vary w.r.t different intra-batch and sequence aggregation functions. Here, the Nemotron-4 15B model is pruned to the Nemotron-3 8B architecture with no retraining. We notice that there is significant variation in zero-shot performance based on the aggregation metric, indicating the importance of selecting the right one. Both (batch=L2, seq=mean) and (mean, mean) perform well; in the remainder of the paper, we use (l2, mean) primarily due to its slightly better performance on the 8T dataset. To further evaluate if these relative rankings hold after retraining, we perform a related experiment: we prune the same 15B model to 8B using: (1) the best ((L2, mean) metric, and (2) a poorly performing (L2, L2) metric, and perform retraining on both for 400 steps ($\sim$1.8B tokens). The results of this experiment are shown in Figure 5. From the Figure, we conclude that these rankings continue to hold post-retraining.
@@
Minitron 4B-instruct 53.09
Gemma-2B-IT 41.63
Llama-3-8B-instruct 50.51
Table 9: Evaluation results on BFCL v2.
@@
Combining Depth and Width (Best Practice #4): we perform a simple experiment to compare the efficacy of width vs. depth pruning. Using the PPL and BI metrics defined in Section 2.2, we remove the 16 least important layers from the Nemotron 15B model based on both metrics to arrive at two variants of depth pruned models. We also perform neuron, head and embedding channel pruning to target the Nemotron-3 8B model and arrive at the width pruned variant. Finally, we combine depth (remove 4 least important layers) and width pruning to arrive at the fourth variant. We report the results of this experiment in Table 10. We notice that even though the depth-width pruned variant has a lower loss post-pruning, we see the results flip around 200 steps of retraining (0.8B tokens); Table 1 and Figure 6 further illustrate this point.
@@
Table 10: Comparison of retraining LM loss across different pruning strategies post retraining with 1.8B tokens. We explore depth only, width only, and a combination of both. Width only strategy though with the least parameter count outperforms the rest.
@@ 4.3 Retraining and Search
Distillation vs. Conventional Training (Best Practice #5): in this experiment, we compare: (1) training a 4B model with random initialization (4B-Random-Init), (2) pruning 15B to 4B, followed by retraining with conventional training (4B-Pruned), and (3) pruning 15B to 4B, and then retraining with distillation using the 15B model as the teacher (4B-Pruned-Distill). Since distillation adds training overheads (additional forward pass on the teacher model), we compare approaches under iso-compute settings. Table 11 shows the results. Here, we observe a significant improvement in MMLU for (3), while both (1) and (2) score randomly. On HellaSwag, (3) > (2) > (1). This clearly demonstrates the superiority of distillation over conventional training after pruning.
@@
Table 11: Accuracy comparison across different strategies to train a 4B model. Pruning the 15B model and distillation results in a gain of 4.8% on Hellaswag and 13.5% on MMLU compared to training from scratch with equivalent compute. Pruning an 8B model instead of a 15B model results in an additional gain of 1% and 4.6% on the benchmarks. Indicates settings with iso-compute.
Choice of Loss Function (Best Practice #5): we experiment with Kullback-Leibler divergence (KLD), MSE, cosine similarity and reverse KLD (R-KLD) to compute $L_{logits}$. Recent work has shown R-KLD [15, 27] to be a better fit than KLD in the SFT/instruction-following setting, and Agarwal et al. [2] claim the choice of loss is task-dependent. We observe from Table 16 and 16 that KLD is the best choice for pruned base model training.
Choice of Losses (Best Practices #6 and #7): typically, a weighted combination of $L_{CLM}$ and $L_{logits}$ is used for distillation. We find that using $L_{logits}$ alone results in the best performance as shown in Table 16. For $L_{is}=L_{emb}+L_{att}+L_{i}+L_{o}$ , we make several observations similar to Lu et al. [31]; these are listed in Appendix A.4 and in Table 17. Most notably, we observe no improvements from using $L_{is}$ when retraining models that don’t prune the depth axis significantly, such as Minitron 8B and Minitron 4B and hence use $L_{logits}$ alone in such cases (see Table 18).
One-shot vs Iterative Pruning and Distillation Across Model Sizes (Best Practice #8): compressing Nemotron-4 15B to Minitron 4B requires an aggressive 73.3% reduction of original model weights. We hypothesize that aggressive one-shot pruning loses out on important capabilities of the base LLM. We thus explore a simple iterative two-step pruning and retraining strategy where we first prune and retrain Nemotron-4 15B to create Minitron 8B ($\sim$46reduction) and further prune and retrain the latter to Minitron 4B ($\sim$50reduction). Table 11 (last two rows) shows the comparison between single-shot and iterative pruning, and demonstrates that iterative achieves a 12% improvement in the MMLU scores compared to the one-shot strategy. During the final retraining step, we observe that using Nemotron-4 15B as the teacher achieves superior results compared to using Minitron 8B. We provide additional ablations on one-shot vs. iterative pruning in Appendix A.5.
Search with Retraining (Best Practice #9): for lightweight neural architecture search, we use the search spaces defined in Table 12 for Minitron 8B and 4B. We further specify a target parameter range of 8 and 4 billion parameters for the respective models, with a tolerance of 5%. With these settings, we obtain 15 and 18 feasible candidates for the 8B and 4B parameter targets, respectively. The architecture configurations for these candidates are provided in Table 19. As described in Section 2.3, we perform lightweight retraining of all feasible candidates. Figure 9 illustrates how validation loss changes for the 8B candidates as training progresses. We notice that relative rankings undergo significant changes up to $\sim$ 300 steps, and then stabilize.